Calculate the density of radon at 325 K and 1.10 atm of pressure.
Answer #1

Ideal gas equation is

PV = nRT

where,

P = Pressure, 1.10atm

V = volume  

n = number of moles

R = gas constant , 0.082057 ( L atm / mol K)

T = Temperature , 325K

n/V = P/RT

n/V = 1.10atm / (0.082057(L.atm/mol K) × 325K)

n/V = 0.04125 mol/L  

Therfore,

Molarity of radon at 325K and 1.10atm = 0.04125 M